Grade 12th passAlgebra

Write a Pythagorean triplet whose one member is. 18

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

2m=18⇒m=18/2=9

m2–1=92–1=81–1=80

m2+1=92+1=81+1=82
∴ (18, 80, 82) is a Pythagorean triplet.

Thanks